# Cross-Modal Few-Shot Learning: a Generative Transfer Learning Framework

### v1

@@ 3.4 Training and Inference
The operational framework of GTL is depicted in Figure 3, details the training and adaptation process for CFSL scenarios. This section outlines the workflow, from initial training on the base dataset to subsequent adaptation for novel data, clarifying the methodologies employed in each phase.
@@
The initial training phase is fundamental as it establishes the distinction between transferable and non-transferable components within the model. During this phase, all modules are trained using the base dataset $D_{\text{base}}$. Specifically, the posterior estimator ($\alpha$), the disturbance estimator and aggregator ($\eta$ and $\delta$), and the generator ($\theta$) are jointly trained using Eq. 12. Afterward, the classifier ($\phi$) is trained using Eq. 13. The primary goal is to robustly encode domain-specific variations and content-specific features into separate latent spaces.
@@
In this phase, the model is exposed to a novel data support set $D_{\text{support}}$. Given our assumption that the relationship between latent representations and visual content remains consistent across both base and novel datasets, we freeze the generator ($\theta$) in its trained state. This decision ensures that the foundational decoding process, which reconstructs visual content from latent representations, remains stable and unaffected by new data variability. We first fine-tune the posterior estimator ($\alpha$), the disturbance estimator, and aggregator ($\eta$ and $\delta$) using Eq. 12, and then update the classifier ($\phi$) with a few labeled examples using Eq. 13.
@@ 4 Experiments
The experiments section consists of Experimental Setup and Benchmark Results. The setup covers datasets, split strategy, evaluation protocol, and implementation details. The results include a summary of outcomes, an ablation study, and a discussion to verify our hypotheses.
4.1 Experimental Setup
@@
We conduct experiments on multiple cross-modal datasets, including two object multi-modal datasets and two biometric multi-modal datasets. Given that sketch is the most common and accessible visual modality beyond RGB, we use multi-modal datasets consisting of both RGB and sketch data to verify the proposed method. For the object multi-modal datasets, Sketchy (Sangkloy et al., 2016) contains 75,471 sketches and 73,002 images across 125 categories, while TU-Berlin (Eitz et al., 2012; Zhang et al., 2016) includes 20,000 sketches and 204,489 images spanning 250 categories. For the more challenging biometric multi-modal datasets which exhibit less inter-class variance, Market-Sketch-1K (Mask1K) (Lin et al., 2023a) contains over 4,700 sketches representing 996 identities in 6 styles, paired with 20,480 matching photographs from Market1501 (Zheng et al., 2015). Additionally, we use SKSF-A (Yun et al., 2024), consists of 938 face-sketch pairs of 134 identities across 7 distinct styles.
@@
For experiments on Sketchy, we follow the dataset splitting scheme used by Bhunia et al. (2022), which contains 64 base classes, while the remaining 61 classes are used as the novel set to make the task more challenging. For TU-Berlin, we split into 125 base and 125 novel classes following Bandyopadhyay et al. (2024). For the smaller biometric multi-modal dataset, all data in Mask1K and SKSF-A are considered novel; we use commonly adopted person and face datasets MSMT17 (Wei et al., 2018) and CelebA-HQ (Karras et al., 2018) as base sets for person and face recognition, respectively. Each base set is further divided into training, validation, and testing subsets (60%:20%:20%) following Bhunia et al. (2022).
@@
We report the results under two settings: all-way-$k$-shot (Ju et al., 2022; Li et al., 2024b) and standard 5-way-$k$-shot (Luo et al., 2023). In the all-way-$k$-shot setting, all classes are presented to the model, with $k$ examples randomly sampled per class to form the support set $S$, while the remaining examples constitute the query set $Q$. In the 5-way-$k$-shot setting, 5 classes are randomly selected per episode, and for each of these 5 classes, $k$ examples are sampled to form the support set $S$, with the remaining examples used for the query set $Q$.
@@
We implement our framework in PyTorch and run experiments on an NVIDIA RTX 2080Ti GPU. ViT-B/16 (Dosovitskiy et al., 2021), pre-trained on the base data with CLIP (Radford et al., 2021) for backbone initialization as the visual encoder to extract the visual representation $\mathbf{x}$. The hyperparameter $d$ is set to 128, and $\lambda$ is set to 1. We use the Adam optimizer Kingma & Ba (2015) for 60 epochs with an initial learning rate of $1e^{-3}$ during the representation learning stage and $1e^{-4}$ during the classification stage. The learning rate decreases to 10% of the original value after 30 epochs, and weight decay is fixed at $1e{-4}$ for all settings.
@@ 4.2 Benchmark Results
We benchmark our method against several few-shot learning methods, including C2-Net (Ma et al., 2024) and ICML23 (Luo et al., 2023), as well as fine-grained retrieval models AGW (Ye et al., 2021), TransReID (He et al., 2021), and CLIP-ReID (Li et al., 2023b).
Methods all-way-1-shot all-way-5-shot 5-way-1-shot 5-way-5-shot
Acc@avg Acc@avg Acc@avg Acc@avg
ICML23 27.6 (29.9 / 25.1) 33.8 (18.1 / 50.6) 28.8 (24.9 / 33.1) 33.7 (16.8 / 52.4)
C2-Net 40.2 (24.7 / 56.7) 42.0 (26.5 / 58.6) 44.2 (40.8 / 48.0) 57.8 (50.1 / 66.3)
AGW 21.1 (23.9 / 18.2) 46.0 (50.4 / 41.4) 45.2 (47.0 / 43.2) 71.7 (82.8 / 59.1)
TransReID 46.0 (23.4 / 70.2) 64.3 (59.9 / 67.0) 78.6 (64.6 / 94.0) 89.6 (83.4 / 96.4)
CLIP-ReID 60.7 (56.2 / 65.4) 81.4 (77.9 / 85.0) 83.2 (76.9 / 90.1) 94.0 (95.3 / 92.5)
Ours 63.8 (58.8 / 69.1) 82.9 (79.8 / 86.3) 84.5 (81.3 / 87.9) 94.1 (95.4 / 92.9)
Methods all-way-1-shot all-way-5-shot 5-way-1-shot 5-way-5-shot
Acc@avg Acc@avg Acc@avg Acc@avg
ICML23 22.9 (34.3 / 21.8) 48.8 (31.5 / 50.4) 35.5 (47.4 / 34.2) 50.9 (48.0 / 51.2)
C2-Net 39.1 (28.8 / 40.1) 50.1 (46.9 / 50.4) 60.1 (60.9 / 70.0) 65.3 (74.9 / 74.3)
AGW 14.7 (32.0 / 13.0) 40.1 (64.2 / 37.9) 69.8 (79.5 / 68.8) 75.9 (89.6 / 74.5)
TransReID 38.0 (29.9 / 38.8) 61.1 (68.9 / 60.3) 83.5 (63.3 / 85.7) 93.3 (92.8 / 93.4)
CLIP-ReID 46.6 (44.2 / 46.8) 74.7 (75.8 / 74.6) 91.5 (79.7 / 92.8) 97.1 (94.9 / 97.3)
Ours 47.1 (46.3 / 47.3) 74.8 (76.1 / 74.7) 92.2 (85.1 / 93.0) 98.0 (98.1 / 98.0)
@@
Sketchy: Table 1 presents the quantitative results on the Sketchy dataset. Our method consistently outperforms all five leading benchmarks, with the highest improvement reaching 4.4% in accuracy across different settings. Notably, in the all-way-1-shot setting, our approach achieves a 3.1% improvement in overall accuracy compared to the best-performing method. Additionally, our model shows gains in both individual modality classification metrics (sketch and RGB) across all settings.
TU-Berlin: The TU-Berlin dataset, which incorporates more abstract representations of data and has a larger scale than Sketchy, but contains fewer sketches. As shown in Table 2, despite the increased difficulty due to the larger number of classes and more challenging multi-modal samples, the proposed GTL method consistently achieves the best performance across all accuracy metrics. Notably, in the 5-way-$k$-shot setting, our approach delivers a 5.4% improvement in sketch modality accuracy, further demonstrating its robustness in handling cross-modal few-shot recognition tasks.
@@
Mask1K: Figures 4(a) and 4(b) present the experimental results on the Mask1K dataset. The increased number of classes and the scarcity of training samples pose significant challenges for existing methods. As shown in the figures, despite the biometric differences being harder to capture, our proposed GTL method achieved notable accuracy improvements of 2.2% and 5.4% for $k=1$ and $k=2$, respectively. Significant improvements were also observed in individual modality classification metrics.
SKSF-A: Figure 4(c) illustrates the performance on the SKSF-A dataset. In this experiment, we report only the average accuracy for the sketch modality, as there is only one RGB image per class available for fine-tuning. The results demonstrate that, regardless of $k=1$ or $k=2$, all methods achieved relatively high accuracy, indicating clear inter-class separability in the dataset. Notably, our method showed an improvement of 11.5% for $k=1$, highlighting its ability to adapt effectively to limited data while maintaining strong performance.
(a) results on Mask1k
(b) results on Mask1k
(c) results on SKSF-A
4.3 Ablation Studies
To verify the effectiveness of each module in GTL framework, Table 4.3 presents the the ablation study results. The first two rows show the results of removing components: “w/o $\mathbf{z}$” excludes all latent variables, using only the classifier ($\phi$) for label prediction; ”w/o $\mathbf{z_{m}}$” removes only the disturbance latent variable $\mathbf{z_{m}}$. The lower rows compare training strategies. $\text{GTL}_{T}$ trains all modules from scratch, while $\text{GTL}_{FT}$ fine-tunes the generator during adaptation instead of keeping it fixed. The last row represents our complete GTL framework, achieving the best results in both settings, as indicated by the bolded accuracy scores.
Table 3: Ablation studies of each component in GTL on Sketchy dataset.
@@ 4.3 Ablation Studies
Comparing the results in Table 4.3, incorporating latent concept learning significantly improves performance. Including $\mathbf{z_{m}}$ (third row) increases RGB accuracy by 18.8% and sketch accuracy by 11.1% over the baseline (first row). Excluding modality-specific variables (”w/o $\mathbf{z_{m}}$”, second row) improves RGB accuracy by 9% and sketch accuracy by 1.7%. Fixing the generator after pre-training (final row) results in the highest average accuracy of 19.6%, outperforming the variant where the generator is fine-tuned. Additional information regarding selecting the hyperparameter of the learnable latent domain number $d$ is provided in Appendix B.
Figure 5: Visualizations of the pair plot of the learned latent representation and the visual representation on Sketchy (left) and Mask1k (right) datasets. Blue stands for model trained on based data, and red denotes the model trained the novel data. The dashed circle with the corresponding color denotes the approximate scope of distributions. The (mix) denotes the trained data are multi-modal, and (RGB) and (Sketch) denote the unimodal data are used for training.
@@ 4.4 Discussions
To validate the invariance of the relationship between latent concepts and visual content and the estimability of in-modality disturbance, we analyzed data distributions. We trained models on base unimodal data, novel multimodal mixed data, RGB data, and sketch data. As shown in Figure 5, pair plots of the generator’s outputs depict the relationship between latent variables $\mathbf{\hat{z}}$ and visual representations $\mathbf{x}$. Significant overlap between base data and novel data distributions supports our assumption. Notably, the overlap with sketch data is greater than with RGB data from the novel set, likely due to differences in underlying data distributions and content diversity. Sketches capture fundamental structures common to the base data, leading to closer alignment in the latent space.
Assumption validation: Estimating in-modality disturbance
To further assess our assumption about estimating in-modality disturbances, we performed t-SNE clustering of the latent representations $\mathbf{\hat{u}}$ from models trained with and without the disturbance estimator. As shown in Figures 6, without the disturbance estimator, latent representations from different modalities—such as RGB images and sketches—are intermixed in the latent space, indicating the model struggles to differentiate modality-specific features. In contrast, when the disturbance estimator is included, t-SNE visualizations reveal distinct clusters for each modality. This demonstrates that the estimator effectively separates modality-specific disturbances and preserves unique characteristics within the latent space.

### v2

@@ 3.4 Training and Inference
This section outlines the workflow, from initial training on the base dataset to subsequent adaptation for novel data, clarifying the methodologies employed in each phase.
@@
The initial training phase is fundamental as it establishes the distinction between transferable and non-transferable components within the model. During this phase, all modules are trained using the base dataset $D_{\text{base}}$. Specifically, the posterior estimator ($\alpha$), the disturbance estimator and aggregator ($\eta$ and $\delta$), and the generator ($\theta$) are jointly trained using Eq. (12). Afterward, the classifier ($\phi$) is trained using Eq. (13). The primary goal is to robustly encode domain-specific variations and content-specific features into separate latent spaces.
@@
In this phase, the model is exposed to a novel data support set $D_{\text{support}}$. Given our assumption that the relationship between latent representations and visual content remains consistent across both base and novel datasets, we freeze the generator ($\theta$) in its trained state. This decision ensures that the foundational decoding process, which reconstructs visual content from latent representations, remains stable and unaffected by new data variability. We first fine-tune the posterior estimator ($\alpha$), the disturbance estimator, and aggregator ($\eta$ and $\delta$) using Eq. (12), and then update the classifier ($\phi$) with a few labeled examples using Eq. (13).
Methods Venues all-way-1-shot all-way-5-shot 5-way-1-shot 5-way-5-shot
Acc@avg Acc@avg Acc@avg Acc@avg
AGW [72] TPAMI’21 21.1 (23.9 / 18.2) 46.0 (50.4 / 41.4) 45.2 (47.0 / 43.2) 71.7 (82.8 / 59.1)
Closer [40] ICML’23 27.6 (29.9 / 25.1) 33.8 (18.1 / 50.6) 28.8 (24.9 / 33.1) 33.7 (16.8 / 52.4)
Style [13] CVPR’23 32.9 (35.2 / 46.6) 40.2 (40.3 / 40.9) 46.6 (51.4 / 41.3) 70.4 (78.1 / 61.9)
LDP-net [77] CVPR’23 40.7 (39.7 / 41.7) 44.1 (41.3 / 47.1) 43.0 (49.8 / 46.5) 66.7 (76.5 / 55.8)
C2-Net [41] AAAI’24 40.2 (24.7 / 56.7) 42.0 (26.5 / 58.6) 44.2 (40.8 / 48.0) 57.8 (50.1 / 66.3)
TransReID [17] ICCV’21 46.0 (23.4 / 70.2) 64.3 (59.9 / 67.0) 78.6 (64.6 / 94.0) 89.6 (83.4 / 96.4)
CLIP-ReID [32] AAAI’23 60.7 (56.2 / 65.4) 81.4 (77.9 / 85.0) 83.2 (76.9 / 90.1) 94.0 (95.3 / 92.5)
Ours - 63.8 (58.8 / 69.1) 82.9 (79.8 / 86.3) 84.5 (81.3 / 87.9) 94.1 (95.4 / 92.9)
Methods Venues all-way-1-shot all-way-5-shot 5-way-1-shot 5-way-5-shot
Acc@avg Acc@avg Acc@avg Acc@avg
AGW [72] TPAMI’21 14.7 (32.0 / 13.0) 40.1 (64.2 / 37.9) 69.8 (79.5 / 68.8) 75.9 (89.6 / 74.5)
Closer [40] ICML’23 22.9 (34.3 / 21.8) 48.8 (31.5 / 50.4) 35.5 (47.4 / 34.2) 50.9 (48.0 / 51.2)
Style [13] CVPR’23 35.5 (37.1 / 35.3) 49.9(50.8 / 49.8) 70.3 (69.9 / 70.3) 76.2 (81.3 / 75.6)
LDP-net [77] CVPR’23 38.7 (39.8 / 38.5) 47.2(52.6 / 46.7) 70.1 (69.7 / 70.2) 74.0 (80.2 / 73.5)
C2-Net [41] AAAI’24 39.1 (28.8 / 40.1) 50.1 (46.9 / 50.4) 60.1 (60.9 / 70.0) 65.3 (74.9 / 74.3)
TransReID [17] ICCV’21 38.0 (29.9 / 38.8) 61.1 (68.9 / 60.3) 83.5 (63.3 / 85.7) 93.3 (92.8 / 93.4)
CLIP-ReID [32] AAAI’23 46.6 (44.2 / 46.8) 74.7 (75.8 / 74.6) 91.5 (79.7 / 92.8) 97.1 (94.9 / 97.3)
Ours - 47.1 (46.3 / 47.3) 74.8 (76.1 / 74.7) 92.2 (85.1 / 93.0) 98.0 (98.1 / 98.0)
@@ 4 Experiments
This section covers experimental setups and benchmark results. The setups detail datasets, split strategy, evaluation protocol, and implementation details, while the results present outcomes, ablation studies, and hypothesis verifications.
4.1 Experimental Setups
@@
We conduct extensive experiments on seven multi-modal datasets: four RGB-sketch, including Sketchy [51] (75K sketches, 73K RGB, 125 classes), TU-Berlin [11, 74] (20K sketches, 204K RGB, 250 classes), Mask1K [35, 76] (4.7K sketches, 996 ids, 6 styles), and SKSF-A [73] (938 pairs, 134 ids, 7 styles); one RGB-infrared SYSUMM01 [65] (22K RGB, 16K infrared, 491 ids); and two RGB-depth, which includes RobotPKU [38] (16K pairs, 180 ids) and Washington RGB-D [27] (250K images, 51 classes).
@@
For experiments on Sketchy, we follow the dataset splitting scheme used by Bhunia et al. [3], which contains 64 base classes, while the remaining 61 classes are used as the novel set to make the task more challenging. For TU-Berlin, we split into 125 base and 125 novel classes following Bandyopadhyay et al. [2]. Similarly, for Washington RGB-D, we split into 25 base and 26 novel classes. For the biometric multi-modal dataset, all data in Mask1K, SKSF-A, SYSUMM01, RobotPKU are considered novel; we use commonly adopted person and face datasets MSMT17 [63] and CelebA-HQ [22] as base sets for person and face recognition, respectively. Each base set is further divided into training, validation, and testing subsets (60%:20%:20%) following Bhunia et al. [3].
@@
We report the results under two settings: all-way-$k$-shot [21, 34] and standard 5-way-$k$-shot [40]. In the all-way-$k$-shot setting, all classes are presented to the model, with $k$ examples randomly sampled per class to form the support set $S$, while the remaining examples constitute the query set $Q$. In the 5-way-$k$-shot setting, 5 classes are randomly selected per episode, and for each of these 5 classes, $k$ examples are sampled to form the support set $S$, with the remaining examples used for the query set $Q$.
@@
We implement our framework in PyTorch and run experiments on an NVIDIA RTX 2080Ti GPU. ViT-B/16 [9], pre-trained on the base data with CLIP [50] for backbone initialization as the visual encoder to extract the visual representation $\mathbf{x}$. The hyperparameter $d$ is set to 128, and $\lambda$ is set to 1. We use the Adam optimizer [25] for 60 epochs with an initial learning rate of $1e^{-3}$ during the representation learning stage and $1e^{-4}$ during the classification stage. The learning rate decreases to 10% of the original value after 30 epochs, and weight decay is fixed at $1e{-4}$ for all settings.
@@ 4.2 Benchmark Results
We benchmark our method against several FSL [41, 40], CD-FSL [13, 77] and fine-grained retrieval models [72, 17, 32].
@@
Sketchy:  Table 2 presents our method consistently outperforms all five leading benchmarks, with the highest improvement reaching 4.4% in accuracy across different settings. Notably, in the aw1s setting, our approach achieves a 3.1% improvement in overall accuracy compared to the second-best method. Additionally, our model shows gains in both individual modality classification metrics (sketch and RGB) across all settings.
TU-Berlin: Table 3 shows despite its larger scale and more abstract samples, GTL achieves the best performance across all accuracy metrics. In the 5-way-$k$-shot setting, it improves sketch modality accuracy by 5.4%, highlighting its robustness in cross-modal few-shot recognition.
Method Washington RGBD RobotPKU
aw1s aw5s 5w1s 5w5s aw1s aw5s 5w1s 5w5s
Closer [40] 25.7 46.7 57.0 68.5 27.5 51.5 54.3 66.0
Style [13] 32.5 53.3 57.8 73.9 38.4 60.9 57.7 75.7
LDP-net [77] 33.6 50.3 60.4 76.2 37.2 59.1 60.0 74.6
CLIP-ReID [32] 39.0 61.2 76.5 81.3 45.4 76.3 81.0 88.2
GTL(ours) 44.4 65.3 80.7 87.6 56.5 82.1 85.9 90.7
Table 4: Quantitative results under Acc@avg of ours and other sota competitors on Washington RGBD and RobotPKU datasets. “aw1s, aw5s, 5w1s, 5w5s” is short for “all-way-1-shot, all-way-5-shot, 5-way-1-shot, 5-way-5-shot”, respectively.
Washington RGB-D: The left of Table 4 supports that the proposed GTL consistently outperforms existing methods, achieving 87.6% accuracy in the 5w5s setting. This demonstrates its strong adaptability in RGB-depth scenarios.
@@
RobotPKU: The right of Table 4 demonstrates GTL achieves sota compared to existing methods and achieves 90.7% in 5w5s, showing its effectiveness in handling cross-modal identity recognition in dynamic environments. The higher performance on both depth datasets is attributed to their limited scene variation, making cross-modal alignment more effective.
Methods SYSUMM01
aw1s aw5s 5w1s 5w1s
AGW [72] 60.5 73.2 63.8 77.9
TransReID [17] 59.2 74.3 66.5 83.3
CLIP-ReID [32] 68.5 81.7 70.4 87.4
Ours 72.1 86.9 75.7 90.0
Table 5: Comparing ours with other competitors on SYSUMM01.
SYSUMM01: As shown in Table 5, our method achieves 72.1% in the aw1s setting and 86.9% in the aw5s setting. For 5-way settings, it reaches 75.7% in 5w1s and 90.0% in 5w5s, demonstrating strong performance across different retrieval configurations.
@@
Mask1K: The high number of classes and limited training samples introduce challenges for recognition. As shown in Figure 5a and 5b, our method achieves accuracy improvements of 2.2% (k=1) and 5.4% (k=2), with notable gains in individual modality classification metrics.
SKSF-A: Since only one RGB image per class is available for fine-tuning, we report only sketch modality accuracy. As illustrated in Figure 5c, all methods achieve relatively high accuracy, indicating clear inter-class separability. Our method improves accuracy by 11.5% for $k=1$, demonstrating strong adaptability to limited data.
@@
Table 6: Ablation studies of each component in GTL on Sketchy.
4.3 Ablation Studies
To verify the effectiveness of each module in GTL framework, Table 6 presents the ablation study results. The first two rows show the results of removing components: “w/o $\mathbf{z}$” excludes all latent variables, using only the classifier ($\phi$) for label prediction; “w/o $\mathbf{z_{m}}$” removes only the disturbance latent variable $\mathbf{z_{m}}$. The lower rows compare training strategies. $\text{GTL}_{T}$ trains all modules from scratch, while $\text{GTL}_{FT}$ fine-tunes the generator during adaptation instead of keeping it fixed. The last row represents our complete GTL framework, achieving the best results in both settings, as indicated by the bolded accuracy scores.
Comparing the results in Table 6, incorporating latent concept learning significantly improves performance. Including $\mathbf{z_{m}}$ (third row) increases RGB accuracy by 18.8% and sketch accuracy by 11.1% over the baseline (first row). Excluding modality-specific variables (“w/o $\mathbf{z_{m}}$”, second row) improves the accuracy of RGB and sketch by 9% and 1.7%. Fixing the generator after pre-training (final row) results in the highest average accuracy of 19.6%, outperforming the variant where the generator is fine-tuned. Additional experiments on the hyperparameter selection of the learnable latent domain number $d$ are provided in the Appendix B.
Figure 6: Distribution analysis among input and generated data.
@@ 4.4 Discussions
Assumption validation: Estimating generative structure
To evaluate the generative structure, we analyze the distributions of input data, decoder-generated output in Figure 6. The left and middle histograms represent the distributions of the original input and the generated output, respectively, while the right plot overlays both distributions. The high similarity between the input and generated data suggests that the decoder effectively reconstructs the original distribution, supporting the validity of our generative assumption.
Assumption validation: Estimating in-modality disturbance
To examine the role of in-modality disturbances, we conduct t-SNE clustering on latent representations $\mathbf{\hat{u}}$ from models trained with and without a disturbance estimator. As shown in Figure 7, without the estimator, latent representations from different modalities (e.g., RGB and sketch) are mixed, indicating the model struggles to distinguish modality-specific features. In contrast, with the estimator, modality-specific clusters emerge, demonstrating its effectiveness in preserving unique characteristics and separating in-modality disturbances.
Figure 8: Visualizations of the pair plot of the learned latent representation and the visual representation on Sketchy (left) and Mask1k (right) datasets. Blue and red stands for model trained on based and novel data, respectively. The dashed circle with the corresponding color denotes the approximate scope of distributions. The (mix) denotes the trained data are multi-modal, and (RGB) and (Sketch) denote the unimodal data are used for training.
@@
To assess the transferability of latent concepts, we trained models on base unimodal data, novel multimodal mixed data, RGB data, and sketch data, and then show their pairwise distributions between latent $\mathbf{\hat{z}}$ and visual representations $\mathbf{x}$ in Figure 8. Significant overlap between base data and novel data distributions in different settings supports our assumption.